OpenStudy (anonymous):

The sum of the first 9 terms of an arithmetric sequence is 81. If the sum of the first and third terms is zero, find the first term and the common difference?

sam (.sam.):

AP \[S_9=81...(1)\] \[S_1+S_3=0...(2)\] ------------------------------ \[S_n=\frac{n}{2}[2a+(n-1)d]\] ------------------------------ For (1), \[2a+(9-1)d=81...(3)\] For (2) \[2a+\cancel{(0)d}+2a+(2)d=0...(4)\] ----------------------------- Simultaneous eq for (3) and (4)

OpenStudy (anonymous):

That is what I thought but my teacher in her solutions did it with the Term Formula which makes no sense

OpenStudy (anonymous):

So instead of Sn = n/2(2a + (n-1)d) she used Tn = a + (n-1)d

sam (.sam.):

I think your teacher meant the sum of \(T_1\) and \(T_3\) is 0

OpenStudy (anonymous):

Because it is a difference equation of the form tn+1 = tn+d (r=1)

sam (.sam.):

AP \[S_9=81...(1)\] \[T_1+T_3=0...(2)\] ------------------------------ \[S_n=\frac{n}{2}[2a+(n-1)d]\] ------------------------------ For (1), \[2a+(9-1)d=81...(3)\] For (2) \[a+a+(2)d=0...(4)\] ----------------------------- Simultaneous eq for (3) and (4)

sam (.sam.):

Exactly

OpenStudy (anonymous):

That is the solution she posted with the question

sam (.sam.):

For the second (2) equation it is \(T_1\)+ \(T_3\)=0

sam (.sam.):

Your teacher wrote \(S_1\) and \(S_3\) = 0 which is misleading

OpenStudy (anonymous):

Well that is what the question says. I believe she has just used the wrong formula

sam (.sam.):

Try it out see if you get the answer

OpenStudy (anonymous):

Thanks for the help. I will get back to studying now

sam (.sam.):

Goodluck, ask if you have any doubts :)

OpenStudy (anonymous):

Will do

OpenStudy (anonymous):

The teacher is right because an arithmetic sequence would look like this: {a, a+d, a+2d, a+3d, ... } Hence (a) (1st term) + (a+2d) (3rd term) = 0 a=-d

OpenStudy (anonymous):

Using the formula tn=a+(n-1)d T9 =81 , a=-d , n=9 81= a + (9-1)d 81= (-d) + (9-1)d 81= -d +8d 7d=81 D=81/7 = 11.57 Since a=-d then a=-11.57

OpenStudy (anonymous):

@hello @.Sam.